At birth the probability that a female will survive to age 65 is approx.8/10 the probability that a male will survive to age 65 is 2/3 what is the probability that both the male and female will be alive at age 65.What is the probablilty that only the male will be alive at 65? Find the probabiltiy that only the female will be alive at 65 and find the probability that at least one of the two will be alive at 65? I just don't understand this at all please help homework due soon. Thanks Susan
Answer by edjones(8007)

Prob both survive to 65: 8/10 * 2/3 =16/30=8/15
.
Only the male will be alive: 2/3 * 2/10 = 4/30 = 2/15
.
Only the female survives: 8/10 * 1/3 = 8/30 = 4/15
.
At least one of the 2 will be alive: 8/15 + 4/15 + 2/15 = 14/15
.
Ed